Use a normal approximation to find the probability of the indicated number of voters. In this​ case, assume that 121 eligible voters aged​ 18-24 are randomly selected. Suppose a previous study showed that among eligible voters aged​ 18-24, 22% of them voted. Probability that fewer than 33 voted

The probability that fewer than 33 of 121 eligible voters voted is .
(Round to four decimal places as needed.)
Answer #1

Answer)

N = 121

P = 0.22

First we need to check the conditions of normality that is if n*p and n*(1-p) both are greater than 5 or not

N*p = 26.62

N*(1-p) = 94.38

Both the conditions are met so we can use standard normal z table to estimate the probability

Z = ( x - mean)/s.d

Mean = n*p = 26.62

S.d = √{(n*p*(1-p)} = 4.55670933898

We need to find

P(x<33)

By continuity correction

P(x<32.5)

Z = (32.5 - 26.62)/4.55670933898

Z = 1.29

From z table, P(z<1.29) = 0.9015

Required probability is 0.9015